prove that n.P(n-1,n-1)=p(n,n)
1
Expert's answer
2021-10-29T03:01:28-0400

Solution:

"LHS=n.P(n-1,n-1)
\\=n\times \dfrac{(n-1)!}{\{(n-1)-(n-1)\}!}
\\=n\times \dfrac{(n-1)!}{0!}
\\=n(n-1)!
\\=n!
\\=\dfrac{n!}{0!}
\\=\dfrac{n!}{(n-n)!}
\\=P(n,n)
\\=RHS"